OpenStudy (anonymous):

Please help @ModernPraise97. There were 300 people at a play. Some paid $5.00 for reserved seats; others paid $3.00 for general seats. Total receipts were $1,080. How many tickets of each kind were sold?

OpenStudy (anonymous):

alright snow so we have our two equations R = reserved g = general seats we know that there are 300 people, soooo r + g = 300 so we can derive this equation 5r + 3g = 1,080 make sense so far?

OpenStudy (anonymous):

I think so ,Adam. It is just that so far in this section that I have been doing something a little different. I have been taught so far that in problems like this I am suppose to let x=one of the variables and x-something is suppose to be the other variable.:)

OpenStudy (anonymous):

its so much easier to say: g = 300 - r then substitute for g in the other equation 5r + 3(300 - r) = 1,080 5r + 900 - 3r = 1,080 2r + 900 = 1,080 2r = 180 r = 90 then substitute for the first equation g + 90 = 300 g = 210 so reserved is 90 and general is 210... but I can try and work it out the way your saying... I'm not sure how far I will get though.

OpenStudy (anonymous):

Oh that is ok maybe I can do it this way Hold on just one sec... I will take a look at this. I have done it this way before just not in this section.:)

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

I totally got this answer! It is just that my teachers are usually kind of picky on whether I did it the way the new section had it. Anyway can we do another problem???... did you get my last text???

OpenStudy (anonymous):

we can do another one sure :)

OpenStudy (anonymous):

As far as you sending me a link of you singing you can send it to jaboyce@gmail.com:)

OpenStudy (anonymous):

Ok. here is another... A housewife paid $7.80 for 5 kilograms of peanuts and 3 kilograms of walnuts. Walnuts are $0.20 a kilogram more than peanuts. How much did she pay per kilogram for each?

OpenStudy (anonymous):

hmmm... I don't know where to start with this... let me examine it..

OpenStudy (anonymous):

Ok:) No problem thanks for helping and take your time:)

OpenStudy (anonymous):

no problem! :) I think it has something to do with... 5p + 3w = 7.80...

OpenStudy (anonymous):

Maybe your right. Let me think about that one quick second.:)

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

I just have one thought on it though... Cold it be that: x=amount for peanuts x+0.20x=amount for walnuts 5x+3(x+0.20x)=$7.80

OpenStudy (anonymous):

yes that looks very good :)

OpenStudy (anonymous):

Then x would =$0.90 that would mean that 5 kilograms of peanuts is $4.50 and... $1.08 amount for walnuts. I guess that wouldn't quit work what do you think??? Could I just be doing some simple math wrong?

OpenStudy (anonymous):

its not 0.20x just 0.20 so: 5x + 3x + 0.60 = 7.80 8x = 7.20 x = 0.90 now we say 4.50 for peanuts and: x = 0.90 + 0.20 x = 1.10 (1.10)3 3.30 4.50 + 3.30 = 7.80 :))

OpenStudy (anonymous):

Ohhh I see! Could I transfer this problem and then work out some more problems with you? I think it just helps to have you right there with me:)

OpenStudy (anonymous):

no prob snow :)

OpenStudy (anonymous):

Alright are you ready, Adam?

OpenStudy (anonymous):

yes ma'am!

OpenStudy (anonymous):

Ok then here it is... Ace's father has 20%-antifreeze solution. How many liters of antifreeze must be added to make a 50%-antifreeze solution? By the way I love my nick name:)Adam

OpenStudy (anonymous):

haha good I'm glad :) thats all the information in the question?!

OpenStudy (anonymous):

The booklet has this formula written out: x=number of liters of antifreeze to be added. 20%(10L) + 100%(x) = 50%(10L +x)

OpenStudy (anonymous):

woah.

OpenStudy (anonymous):

It also has this: (Hint: the product to be added is 100% antifreeze)

OpenStudy (anonymous):

It also tells me to make all the percents into decimal fractions if that helps:)

OpenStudy (anonymous):

ok... 20 50 ---(10) + x = --- (10 + x) 100 100

OpenStudy (anonymous):

sorry I'm not really seeing what you did here. What are the---

OpenStudy (anonymous):

Oh sorry I am just being stupid I see what u did know:)

OpenStudy (anonymous):

I mean "now" lol

OpenStudy (anonymous):

but would it not be: 0.20(10L) +0.100x=0.50(10L+x)

OpenStudy (anonymous):

yeah... but its like... what now? lol

OpenStudy (anonymous):

I think I just solve for x, but then there is that L?!lol

OpenStudy (anonymous):

This is what I have so far: 0.20(10L) +0.100x=0.50(10L+x) 2L+0.100x=5L+0.50x 0.50x=3L Like you said though it's like what now??? lol

OpenStudy (anonymous):

hahaha yeah.. thats weird...

OpenStudy (anonymous):

Do you have any ideas?:) Your suppose to be the smart one here junior!:)

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

haha very true but I actually have to go :O im sorry..

OpenStudy (anonymous):

text me though... ok snow? :)

OpenStudy (anonymous):

Don't worry I will, Adam! Do you mean you can't use the computor anymore or you won't be available?

OpenStudy (anonymous):

Iwill be on the move, but will still have my phone :)

OpenStudy (anonymous):

bye!

OpenStudy (anonymous):

Ok sounds great I will text you. I just have to eat supper in about 10 minutes. So be on the watch for my text in about 30 to 40 minutes, OK? Bye Adam

OpenStudy (anonymous):

alright :) bye snow